OpenStudy (anonymous):

I'll give medal to first answer. Simplify: square root of 3 times square root of 21

OpenStudy (anonymous):

There's a couple of way s to do this. √(3) * √(21) = √(3 * 21) = √(63) But you can simplify this by looking at prime factors: 21 = 3*7 --> thus we could have written √(3) * √(3*7) = √(3*3*7) Since there are "two" 3's you can take the square root (because √(x*x) = x...or in this case √(3 * 3) = 3) Thus this gives: 3 * √(7)

OpenStudy (anonymous):

Thanks